# Wonder Wins Ways: Curiosity-Driven Exploration through Multi-Agent Contextual Calibration

### v1

@@ 4.4 Intrinsic Reward Module
While the intrinsic reward may be challenging due to the non-Bayesian parameterization of the Cermic model, We emperically approximate it with an IB-trained representation and a Gaussian marginal, yielding a tractable lower-bound bonus that supports robust exploration in noisy environments (see Appendix D).
@@ 5.1 Task Setup
Figure 2: Visualization of agent observation embeddings $s_{t}$ and latent states $x_{t}$. From VMAS: Balanc Task. (Left) Comparative $s_{t}$ distributions for agents under Cermic-augmented vs. baseline algorithms. (Right) Influence of curiosity-driven latent states $x_{t}$ on task exploration.
@@ 5.2 Comparison with State-of-the-Art
Additionally, Agents driven by single-agent curiosity formulations (e.g., MAPPO-DB), which ignore peers’ intentions, are persistently distracted by others’ behaviors even in later training. Our contextual calibration explicitly addresses this issue by transforming socially induced randomness into a learnable signal. To further validate this, we evaluate how per-agent contributions change with the number of agents in the Balance* task. Results show that MAPPO-DB’s per-agent contribution decreases as agent count increases, while Cermic mitigates this degradation:
@@ 5.3 Qualitative and Quantitative Analysis
Figure 3: Learning curves on Balanc.
@@
Quantitative Analysis. Fig.˜3 illustrates the mean episodic return curves on the VMAS: Balanc task, with comprehensive results across additional tasks deferred to Appendix F. While Cermic-augmented algorithms ultimately achieve superior final performance, their learning curves can exhibit a slower initial ramp-up compared to some baselines. We attribute this initial lower sample efficiency to the inherent challenge of learning effective multi-agent intention modeling from scratch, a difficulty induced by severe partial observability.
@@ 5.4 Ablation Study
Loss Module. We prove the influence of each component in Cermic’s composite loss function, with results presented in Fig.˜4. Our findings demonstrate that each loss component is crucial for achieving optimal performance: specifically, removing the exploration loss ($\mathcal{L}_{\text{explore}}$) significantly hindered adaptation to the sparse-reward environment, as evidenced by slower learning progress in the corresponding curve; conversely, omitting the exploitation loss ($\mathcal{L}^{\rm UB}_{\text{exploit}}+\mathcal{L}^{\rm LB}_{\text{exploit}}$) led to an inability to effectively filter noise, causing notable instability and fluctuations throughout learning.
Hyperparameter $\epsilon$. We investigated the impact of varying the chance constraint tolerance $\epsilon$ on Cermic’s behavior, with results in Fig.˜5. Our findings reveal that smaller values of $\epsilon$ induce more conservative exploration strategies. The underlying reason is that a smaller $\epsilon$ compels the agent to favor actions that keep $\Psi_{t}$ reliably low, thereby prioritizing exploitation and reducing the pursuit of novelty. Consequently, while diminished exploration due to smaller $\epsilon$ leads to lower average task performance, it concurrently results in reduced variance across multiple seeds.
Figure 5: Latent state visualization: Balanc.
More Ablations. We conducted additional ablations on the intention model architecture and a key hyperparameter $\alpha$ (Table˜3). Employing more sophisticated intention models consistently improved task performance, underscoring the benefit of richer contextual feature $f_{n}^{o}$ extraction for Cermic’s calibration. Regarding $\alpha$, which scales our calibrated intrinsic curiosity, larger values typically reduced immediate task returns due to increased exploration, but concurrently enhanced learning stability and reduced performance variance, highlighting a trade-off between rapid exploitation and robust discovery.
@@ Appendix B Derivation with Cantelli Inequality
where $\mathcal{P}_{\rm cal}$ refers to the set of distributions under multi-agent contextual intention uncertainty in the main paper. To this end, Eq.˜18 formulates the problem as a bi-level optimization. The outer layer finds the worst-case mean $\mu_{1}$ and variance $\sigma_{1}^{2}$ within the ambiguity set $S$. The inner layer finds the worst-case probability $\mathbb{P}_{r}(\tilde{s}\leq b)$ for a given mean and variance.
@@
In practical multi-agent reinforcement learning scenarios, the desired confidence level $1-\epsilon$ is positive. For the chance constraint $\inf\mathbb{P}_{r}(\tilde{s}\leq b)\geq 1-\epsilon$ to hold with $1-\epsilon>0$, we must be in the regime where $b\geq\mu_{1}$. If $b<\mu_{1}$, the infimum probability would be 0, violating the constraint. Therefore, for any relevant $(\mu_{1},\sigma_{1})\in S$, the condition $b\geq\mu_{1}$ must be satisfied. This implies $b\geq\sup_{(\mu_{1},\sigma_{1})\in S}\mu_{1}=\sqrt{\gamma_{1}\Sigma_{h}}$ (assuming $\Sigma_{h}>0$). Consequently, Eq.˜18 simplifies to:
@@
Solving the optimization problem in Eq.˜20 over the set $S$ (defined by moment bounds $|\mu_{1}|\leq\sqrt{\gamma_{1}\Sigma_{h}}$ and $\mu_{1}^{2}+\sigma_{1}^{2}\leq\gamma_{2}\Sigma_{h}$) yields the following worst-case probability:
@@ D.2 Connection to LSVI-UCB Bonus
D.3 Empirical Implementation
The intrinsic reward $r^{i}$ relies on a Bayesian view of model parameters $\Theta$. As our Cermic model uses non-Bayesian neural networks, this section outlines a practical approximation for $r^{i}$.
@@ Appendix F Supplementary Experimental Results
Performance Curve Comparison. We present further comparative experiments showcasing the performance of Cermic against other methods in Fig.˜6 below. The results demonstrate that Cermic consistently surpasses prior SoTA models and other curiosity-driven approaches, achieving superior task performance. Furthermore, it is observable that our method exhibits more stable learning curves, characterized by narrower error bands. We attribute this enhanced stability and performance to Cermic’s ability to effectively filter noisy signals while concurrently encouraging robust exploration.
@@
Robustness to Noise. To further validate the noise-filtering capabilities of Cermic, we conducted comparative experiments in environments augmented with “random box” noise, as depicted in Fig.˜7. The results indicate that, in these noisy conditions, Cermic maintains more stable learning curves and narrower error bands compared to other algorithms. Notably, this advantage in stability is more pronounced than in the noise-free task counterparts (Fig.˜6). Furthermore, we observe that the error bands in the Simple Adversary task are smaller than those in the Ball Trajectory task. This difference can be attributed to the richer multi-agent contextual information available in the former, which includes signals from both adversaries and teammates, whereas the latter primarily involves information from a single partner. Consequently, the increased contextual feedback translates to a stronger noise-filtering effect, as reflected in the learning curves. This underscores Cermic’s ability to leverage multi-agent information to enhance task understanding and mitigate the impact of environmental stochasticity.
Role Diversity and Task Adaptation. While Cermic consistently improves performance, we observe a more substantial performance gain over traditional baselines in tasks involving diverse agent roles (e.g., both cooperation and competition, as in Simple Adversary) compared to purely cooperative/competitive settings (e.g., Ball Trajectory); see Fig.˜7. We attribute this to the challenge faced by traditional methods in interpreting heterogeneous agent behaviors under partial observability, where inferring intent and role from observations is inherently difficult. In contrast, Cermic’s task-adaptive weighting mechanism ($\gamma$) enables agents to better infer and adapt to others’ intentions and plans, thereby enhancing performance in complex social interactions.
Incremental Enhancement. We conducted an ablation study to analyze the incremental impact of Cermic’s components on task performance. This involved comparing Mappo, Mappo-Cermic, and Mappo-Cermic with a pre-trained memory module. The results, presented in Fig.˜8, intuitively illustrate the contribution of each component to the baseline algorithm’s performance. It can be observed that the core Cermic module primarily serves to elevate overall performance, while also providing a modest improvement in learning stability. The subsequent integration of a pre-trained memory module further enhances this stability, leading to more consistent learning trajectories.
Figure 8: Incremental impact of Cermic components on VMAS tasks.
Testing Agent Generalization. We further investigated the generalization capabilities of agents by evaluating their performance when trained under one reward density (dense or sparse) and tested under a sparse-reward setting. The results are presented in Table˜4. A consistent trend observed across all algorithms is a performance degradation when agents trained on dense rewards are deployed in sparse-reward scenarios. However, Cermic demonstrates a notable ability to mitigate this performance drop. We attribute this to Cermic’s capability to enable agents to understand the task by observing other agents, rather than solely relying on trial-and-error learning. These findings underscore the importance of research in sparse-reward reinforcement learning and highlight the potential for Cermic’s broad applicability.
Method Balance Give Way Passage
[…]
Cpm 61.0 58.2 -2.8 3.73 3.58 -0.15 162 153 -9
Qplex-Ices 63.2 60.7 -2.5 3.96 3.85 -0.11 164 159 -5
Mappo-Cermic 67.3 65.6 -1.7 3.94 3.82 -0.12 172 168 -4
Visualization of Intrinsic Reward. To illustrate the temporal behavior of Cermic’s intrinsic reward, we visualized its values throughout a complete episode of the Balance task, as shown in Fig.˜9. A clear trend is the gradual decrease of this intrinsic reward over time, indicating a diminishing role of curiosity-driven exploration as the agent gains familiarity with the task environment. Furthermore, peaks in the intrinsic reward consistently coincide with an agent encountering novel situations, such as initial environmental entry or first-time interactions with other agents. This observation aligns with our design objective for the intrinsic reward, which is to incentivize exploration of unfamiliar states and interactions.

### v2

@@ 4.4 Intrinsic Reward Module
While the intrinsic reward may be challenging due to the non-Bayesian parameterization of the Cermic model, We empirically approximate it with an IB-trained representation and a Gaussian marginal, yielding a tractable lower-bound bonus that supports robust exploration in noisy environments as follows:
@@ 5.2 Comparison with State-of-the-Art
Additionally, Agents driven by single-agent curiosity formulations (e.g., MAPPO-DB), which ignore peers’ intentions, are persistently distracted by others’ behaviors even in later training (see Table˜2). Our contextual calibration explicitly addresses this issue by transforming socially induced randomness into a learnable signal. To further validate this, we evaluate how per-agent contributions change with the number of agents in the Balance* task. Results show that MAPPO-DB’s per-agent contribution decreases as agent count increases, while Cermic mitigates this degradation:
@@ 5.3 Qualitative and Quantitative Analysis
Figure 2: Visualization of agent observation embeddings $s_{t}$ and latent states $x_{t}$. (Left) Comparative $s_{t}$ distributions for agents under Cermic-augmented vs. baseline algorithms. (Right) Influence of curiosity-driven latent states $x_{t}$ on task exploration.
Figure 3: The impact of pretrained intention memory module on Cermic’s performances
@@
Quantitative Analysis. Fig.˜3 illustrates the mean episodic return curves on the VMAS task, with comprehensive results across additional tasks deferred to Appendix F. While Cermic-augmented algorithms ultimately achieve superior final performance, their learning curves can exhibit a slower initial ramp-up compared to some baselines. We attribute this initial lower sample efficiency to the inherent challenge of learning effective multi-agent intention modeling from scratch, a difficulty induced by severe partial observability.
@@ 5.4 Ablation Study
We conduct three ablations on Cermic, summarized in Table˜3: (i) Loss module: Ablating the exploration loss $\mathcal{L}_{\text{explore}}$ markedly impedes adaptation in sparse-reward settings, yielding slower learning; removing the exploitation losses $\mathcal{L}^{\rm UB}_{\text{exploit}}+\mathcal{L}^{\rm LB}_{\text{exploit}}$ prevents effective noise filtering, leading to instability and pronounced fluctuations. (ii) $\alpha$: increasing the curiosity scale $\alpha$ (for our calibrated intrinsic curiosity) typically lowers immediate returns via greater exploration, yet improves learning stability and reduces performance variance, reflecting an exploration–exploitation trade-off. (iii) Intention model: More expressive intention models that better extract contextual features $f_{n}^{o}$ consistently improve task performance.
@@
5.5 Generalization Across Reward Densities
We further investigated the generalization capabilities of agents by evaluating their performance when trained under one reward density (dense or sparse) and tested under a sparse-reward setting. The results are presented in Table˜4. A consistent trend observed across all algorithms is a performance degradation when agents trained on dense rewards are deployed in sparse-reward scenarios. However, Cermic demonstrates a notable ability to mitigate this performance drop. We attribute this to Cermic’s capability to enable agents to understand the task by observing other agents, rather than solely relying on trial-and-error learning. These findings underscore the importance of research in sparse-reward reinforcement learning and highlight the potential for Cermic’s broad applicability.
Method Balance Give Way Passage
[…]
Cpm 61.0 58.2 -2.8 3.73 3.58 -0.15 162 153 -9
Qplex-Ices 63.2 60.7 -2.5 3.96 3.85 -0.11 164 159 -5
Mappo-Cermic 67.3 65.6 -1.7 3.94 3.82 -0.12 172 168 -4
@@ 6 Conclusion
Acknowledgment
This paper was supported by the Natural Science Foundation of China under Grant 62303307, 62225309, U24A20278, 62361166632, U21A20480, and Sponsored by the Oceanic Interdisciplinary Program of Shanghai Jiao Tong University, and in part by National Key Laboratory of Human Machine Hybrid Augmented Intelligence, Xi’an Jiaotong University (No. HMHAI-202408).
@@ Appendix B Derivation with Cantelli Inequality
where $\mathcal{P}_{\rm cal}$ refers to the set of distributions under multi-agent contextual intention uncertainty in the main paper. To this end, Eq.˜20 formulates the problem as a bi-level optimization. The outer layer finds the worst-case mean $\mu_{1}$ and variance $\sigma_{1}^{2}$ within the ambiguity set $S$. The inner layer finds the worst-case probability $\mathbb{P}_{r}(\tilde{s}\leq b)$ for a given mean and variance.
@@
In practical multi-agent reinforcement learning scenarios, the desired confidence level $1-\epsilon$ is positive. For the chance constraint $\inf\mathbb{P}_{r}(\tilde{s}\leq b)\geq 1-\epsilon$ to hold with $1-\epsilon>0$, we must be in the regime where $b\geq\mu_{1}$. If $b<\mu_{1}$, the infimum probability would be 0, violating the constraint. Therefore, for any relevant $(\mu_{1},\sigma_{1})\in S$, the condition $b\geq\mu_{1}$ must be satisfied. This implies $b\geq\sup_{(\mu_{1},\sigma_{1})\in S}\mu_{1}=\sqrt{\gamma_{1}\Sigma_{h}}$ (assuming $\Sigma_{h}>0$). Consequently, Eq.˜20 simplifies to:
@@
Solving the optimization problem in Eq.˜22 over the set $S$ (defined by moment bounds $|\mu_{1}|\leq\sqrt{\gamma_{1}\Sigma_{h}}$ and $\mu_{1}^{2}+\sigma_{1}^{2}\leq\gamma_{2}\Sigma_{h}$) yields the following worst-case probability:
@@ Appendix F Supplementary Experimental Results
Performance Curve Comparison. We present further comparative experiments showcasing the performance of Cermic against other methods in Fig.˜4 below. The results demonstrate that Cermic consistently surpasses prior SoTA models and other curiosity-driven approaches, achieving superior task performance. Furthermore, it is observable that our method exhibits more stable learning curves, characterized by narrower error bands. We attribute this enhanced stability and performance to Cermic’s ability to effectively filter noisy signals while concurrently encouraging robust exploration.
@@
Robustness to Noise. To further validate the noise-filtering capabilities of Cermic, we conducted comparative experiments in environments augmented with “random box” noise, as depicted in Fig.˜5. The results indicate that, in these noisy conditions, Cermic maintains more stable learning curves and narrower error bands compared to other algorithms. Notably, this advantage in stability is more pronounced than in the noise-free task counterparts (Fig.˜4). Furthermore, we observe that the error bands in the Simple Adversary task are smaller than those in the Ball Trajectory task. This difference can be attributed to the richer multi-agent contextual information available in the former, which includes signals from both adversaries and teammates, whereas the latter primarily involves information from a single partner. Consequently, the increased contextual feedback translates to a stronger noise-filtering effect, as reflected in the learning curves. This underscores Cermic’s ability to leverage multi-agent information to enhance task understanding and mitigate the impact of environmental stochasticity.
@@
Role Diversity and Task Adaptation. While Cermic consistently improves performance, we observe a more substantial performance gain over traditional baselines in tasks involving diverse agent roles (e.g., both cooperation and competition, as in Simple Adversary) compared to purely cooperative/competitive settings (e.g., Ball Trajectory); see Fig.˜5. We attribute this to the challenge faced by traditional methods in interpreting heterogeneous agent behaviors under partial observability, where inferring intent and role from observations is inherently difficult. In contrast, Cermic’s task-adaptive weighting mechanism ($\gamma$) enables agents to better infer and adapt to others’ intentions and plans, thereby enhancing performance in complex social interactions.
Visualization of Intrinsic Reward. To illustrate the temporal behavior of Cermic’s intrinsic reward, we visualized its values throughout a complete episode of the Balance task, as shown in Fig.˜6. A clear trend is the gradual decrease of this intrinsic reward over time, indicating a diminishing role of curiosity-driven exploration as the agent gains familiarity with the task environment. Furthermore, peaks in the intrinsic reward consistently coincide with an agent encountering novel situations, such as initial environmental entry or first-time interactions with other agents. This observation aligns with our design objective for the intrinsic reward, which is to incentivize exploration of unfamiliar states and interactions.